OpenStudy (anonymous):

Choose the correct description of the graph of the inequality x − 12 less than or greater to 5x. A. Open circle on −3, shading to the left. B. Closed circle on −3, shading to the left. C. Open circle on −3, shading to the right. D. Closed circle on −3, shading to the right.

OpenStudy (anonymous):

solve for x first

OpenStudy (anonymous):

Do you know what to do first?

OpenStudy (anonymous):

..well i thought i would add 12 to both sides first but idk

OpenStudy (anonymous):

no you want to add or subtract the x until you get one x on one side

OpenStudy (anonymous):

\[x − 12\le 5x \\x -x − 12\le 5x-x \\ − 12\le 4x \\− 3\le x \]

OpenStudy (anonymous):

From that which one do you think it is?

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

No we want values to be greater than -3 . like 0 and 1 and 5 are those numbers to the left or right of -3? You got the closed circle part right

OpenStudy (anonymous):

ohh to right thank u!!!

OpenStudy (anonymous):

to the right*

OpenStudy (anonymous):

Yup